Intro to Econometric Theory Heinz School, Carnegie Mellon University 90-906, Spring 2005-6

Homework #5, due Monday, May 1, 2006

1. Do problem 10 on page 115 of Greene

2. The Medical Expenditure Panel Survey is an annual survey which collects information about medical expenditures, income, employment, demographics, health information, &c for a representative sample of Americans.

I have prepared an extract of these data for 1996, and it is available on the course website. The following are the columns in the data, in order:

Variable  Meaning
age  age of person in years
sex  sex of person, 1=male & 0=female
income  income in 1996 $
employed  1=employed, 0=not employed
insured  1=had health insurance, 0=not
health perceived health status, higher is sicker
spending  spending on health care, 1996 $

To begin with, let’s consider this model:

spending_{i} = β_{1} + β_{2}income + β_{3}age + β_{4}sex
+ β_{4}employed + β_{5}insured + β_{6}health  (1)

(a) Test this model for heteroskedasticity. Discuss your findings and and what they mean for inference.

(b) How much more do people in very good health status spend than do people in excellent health status (estimate and CI). Make sure your answer is correct in the presence of heteroskedasticity.

(c) Test whether insurance affects spending for people of different health statuses differently and discuss. Make sure your anser is correct in the presence of heteroskedasticity.